A health psychologist tests a new intervention to determine if it can change healthy behaviors among siblings. To conduct the this test using a matched-pairs design, the researcher gives one sibling an intervention, and the other sibling is given a control task without the intervention. The number of healthy behaviors observed in the siblings during a 5-minute observation were then recorded.

Intervention
Yes No
6 4
3 5
6 4
6 5
6 4
5 4

(a) Test whether or not the number of healthy behaviors differ at a 0.05 level of significance. State the value of the test statistic. (Round your answer to three decimal places.)
Incorrect: Your answer is incorrect.

State the decision to retain or reject the null hypothesis.

Retain the null hypothesis.Reject the null hypothesis.    Correct: Your answer is correct.


(b) Compute effect size using eta-squared. (Round your answer to two decimal places.)
Incorrect: Your answer is incorrect.
Answer #1

Given that,
null, H0: Ud = 0
alternate, H1: Ud != 0
level of significance, α = 0.05
from standard normal table, two tailed t α/2 =2.571
since our test is two-tailed
reject Ho, if to < -2.571 OR if to > 2.571
we use Test Statistic
to= d/ (S/√n)
where
value of S^2 = [ ∑ di^2 – ( ∑ di )^2 / n ] / ( n-1 ) )
d = ( Xi-Yi)/n) = 1
We have d = 1
pooled variance = calculate value of Sd= √S^2 = sqrt [ 18-(6^2/6 ] / 5 = 1.549
to = d/ (S/√n) = 1.581
critical Value
the value of |t α| with n-1 = 5 d.f is 2.571
we got |t o| = 1.581 & |t α| =2.571
make Decision
hence Value of |to | < | t α | and here we do not reject Ho
p-value :two tailed ( double the one tail ) - Ha : ( p != 1.5811 ) = 0.1747
hence value of p0.05 < 0.1747,here we do not reject Ho
ANSWERS
---------------
a.
null, H0: Ud = 0
alternate, H1: Ud != 0
test statistic: 1.581
critical value: reject Ho, if to < -2.571 OR if to > 2.571
decision: Do not Reject Ho
p-value: 0.1747
we do not have enough evidence to support the claim that whether or not the number of healthy behaviors differ at a 0.05 level of significance.
b.
effect size = d = ( Xi-Yi)/n) /S.D pooled
S^2 = sqrt [ 18-(6^2/6 ] / 5 = 1.549
S.D pooled = sqrt (1.549) = 1.2445
effect size = d/S.D pooled
effect size = 1/1.2445
effect size = 0.8035
large effect